Summary of the 3mm GMVA Session of September 26-29, 2013 -------------------------------------------------------- Stations participating: Eb, On, PV, PdB, Yb, VLBA(7) plus : KVN (Ky, Ku, Kt) at 3mm : Nt at 7mm : Eb at 7mm from 3h15 UT - 10h50 UT on day 270 An ftp-fringe test showed fringes between the RDBE Stations (Eb, VLBA_PT) but not between the stations which use DBBCs (PV, Yb, On). At present it is unclear if this is a problem with the polyphase filterbank configuration of the DBBC or a problem in the setup of the DifX correlator. Due to bad weather PV could observe only for the first 10-12 hrs. Due to the lack of ftp fringe detection PdB dropped at 00h07 UT on day 270. The weather at Eb, Yb, On was reasonable good. At the VLBA, the station VLBA_LA did not observe due to a technical problem of the telescope (broken AZ axle). VLBA observations were interrupted on each day by 1-2 hrs for USNO PT-MK monitoring. Individual Station Reports: Effelsberg: day 269 - Started in time. Effelsberg partly cloudy, but clearing up. Tsys ~300K T=13.5 Celcius, H=93.5%. We point on 0716+714 and can see it. day 270 - scheduled switch to 7mm, observing part of 3C84 with 7mm secondary focus RX during 3h13 UT - 10h50 UT 11h15 UT back to 3mm VLBI day 270 - 272: clear sky, very good weather some pointing probelms during daytime because of sun Pico Veleta: day 269 - partly cloudy, wind close to the safety limit of the telescope. Tsys about 150 K <= 19UT wrong LO setting discovered, LO off by 2 MHz, reset LO to correct value at 19 UT day 270 - 11h30 UT stop observing due to high wind day 271 - bad weather, no observation day 272 - bad weather, no observation Note added after fringesearch: only LCP, RCP is being overwritten, wrong dbbctrk commands in station procedure. PdBure: day 269 - 14 UT Problems with Field System, FS and Formatter clock drift and diverge 18:25 UT restart FS, now better 19:00 UT weather good, local phases stable day 270 - PdB decided to stop observing, since correlator could not return results from ftp fringe test. Will resume when correlator says that fringes are found. day 271-272: cannot find fringes to PdB, DifX might have a problem with zoom-mode correlation and/or sideband swapping (USB,LSB) PdB observes local projects, no VLBI Yebes: day 269 - 14h30 UT Clear sky, prec water=8.6 mm, opac=0.16. T=29 C, HR=29% day 270-272 good, no obvious problems day 273 - check quarter wave plate, realize that it was not mounted correctly So, the last 4 days the polarization has been RCP. Onsala: We have observed c132a. None of the scripts we use work for the PFB-mode. i Neither does the Tsys calculation in the FS. We lost some time around 05 UTC (day 270) due to the Mark5B+. Summary for mt002 (triggered obs on SgrA*, Europe only): --------------------------------------------------------- mt002a: not observed, waether bad mt002b: not observed, weather bad mt002c: observed Effelsberg, Pico, PdB, Yebes Yebes dropps out earlier due to a scheduled other project individual comments: PV: We have started recording mt002c in time. The weather is good and Tsys about 117 K. PdB: started MT002c on time, Tsys ~90 K, convective clouds but phases ok EB: clear sky, no probelms Yb: first scan lost, rest ok, will stop 17h00 UT for another observation. VLBA: !!! KP may have swapped polarization during Sep 2013 - Feb 2014, check !!!! VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C132A Observer: Krichbaum, T. Project type: VLBI Obs filename: c132acrd.* Date/Day: 2013SEP26/269 Ants Scheduled: NL FD PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Alan Kerr GPS delta, vclock: OK, UT1-UTC: 0.02 1356 Begin 1356-1356 *NL 3mm high Tsys. FE 50 2160 1356-1356 *LA not observing; broken AZ axle (C133781) AMECH 100 4320 1722-1745 *KP not recording. problems switching banks on the REC 100 23 mark5c. 1957-2001 *OV out of point. power glitch. COM 100 4 2046-2048 *OV out of point. power glitch. COM 100 2 2145 New Operator: Paul Padilla 2151-2153 *OV autostowed in wind gust WEA 100 2 0000 Date is 2013SEP27/270. 0400-0445 *PT,MK removed for USNO project. SCH 200 90 0536 Operator is A. Sowinski. 0620-0712 *PT,MK removed for backup USNO project. SCH 200 104 (N3270B) 1051-1052 *PT Autostow, power glitch. COM 100 1 1345 New operator is Alan Kerr 1530-1552 *PT out to check clock settings. TIME 100 22 1605 %BR light rain 2000-2056 *PT, MK out for USNO observation. SCH 200 112 2146 New Operator: Paul Padilla 2236-2244 *PT,MK executor being updated OTHER 100 8 2325-0012 *PT not recording, invesitgating possible reasons REC 100 47 0000 Date is 2013SEP27/270. 0539 Operator is A. Sowinski. 0555-0651 *PT,MK removed for USNO N3271. SCH 200 112 1252-1253 *PT Autostow, repointed. COM 100 1 1339 New operator: T. Perreault 1406 Communications with sites are now down, planned power outage at DSOC. Closing down log, observing will continue to run being sporadically monitored by site techs and operations. 1726 Recovered from power outage. Back to monitoring observation at DSOC 1811 KP wea: clear, no precipitation 0000 Date is 2013SEP29/272. 0140 New operator is Betty Ragan. 0239 %BR windy. 0337 %BR raining. 0352-0400 *NL record rate was zero. Checked the module, REC 100 8 and it looked OK. Restarted DRS, and it appears to be recording OK again. 0611 %NL raining. 0651-0751 *PT,MK removed for USNO observation. SCH 200 120 1329 New operator: T. Perreault 1352 EB observing ends 1356 End. Downtime Summary: Total downtime : 7136 min Percentage downtime of observing: 23.6% Average downtime per hour : 14.2 min Total scheduled observing time (# Antennas): 30240 min (7) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard.